5 The parent peak in the mass spectrum shown below is at 100 80 6o 40 C 20 10 20 3 40 070 90 m/z (a) m/z 58 (b) m/z -71 (c) m/z - 86 (d) m/z 43 6 In mass spectrometry of the following organic compound, which of the following fragmentation pathway is not likely to occur? H3C CH H3C (a) dehydration (b) Alpha-cleavage (c) McLafferty rearrangement (d) All of these
Answer #1

5. The parent peak in the given mass spectrum is at m/z = 86.

In the mass spectrum, the molecular ion (parent) peak or the M+ peak is at m/z = 86. The tallest peak is the base peak. The relative abundance of the ion at m/z = 43 is the maximum hence it is the base peak. There is also one small peak at m/z = 87 which is the M+1 peak that corresponds to a 13C since 1.1% of the carbon atoms that occur in nature are 13C and it being heavier than 12C corresponds to the M+1 peak. There are no more peaks after the M+1 peak because there can be no ion having a greater molecular mass than the molecular (parent) ion explaining the fact that the peak at m/z = 86 is the parent peak or the molecular ion peak.

Hence correct option is c.

6. Fragmentation in the given compound will not occur by any of the given fragmentation methods.

McLafferty rearrangement - It occurs in an organic molecule that contains a keto group. In the compound given, a keto functionality is missing. Hence McLafferty rearrangement does not occur here.

Alpha cleavage -  For alpha cleavage to occur, a compound must contain heteroatoms such as oxygen, sulphur or nitrogen where bonds break adjacent to the heteroatoms that generate a resonance stabilised carbocation. In the compound given, a heteroatom is missing. Hence alpha cleavage does not occur here.

Dehydration - For dehydration to occur i.e for the loss of a water molecule, an alcohol functionality should be present. Here again, a OH functional group is missing. Hence dehydration too does not take place here.

Hence the correct option is d.